Cpuld you help me with naming the point on the line when the point-slope form is: y=7-2(x-3)?
Your help is greatly appreciated!
P.S. I got (-5,7) and it was incorrect.

Answer by richard1234(7193)

Pick some x value and then find the value of y. For example, if x = -5, then y equals 7 - 2(-5 - 3) = 7 - 2(-8) = 7 + 16 = 23, so (-5,23) is one point on the line.